# Error Correction Code Transformer: From Non-Unified to Unified

### v1

@@ 5 Analysis
We delve into the influence of Transformer encoder layers on overall performance, examining both the role of the dimension parameter $d_{k}$ and the efficacy of sparse masked attention mechanisms. Furthermore, we conduct a comparative analysis of the computational complexity inherent in our methodology versus that of alternative approaches.
5.1 Impact of the Transformer Encoder Layers and the Dimensions $d_{k}$
Figure 10 illustrates the CBER performance across various Transformer encoder layers and the dimensions $d_{k}$. It can be observed that the number of Transformer encoder layers exerts a greater impact on performance than the dimension $d_{k}$. This increase in depth effectively enlarges the parameter search space, potentially enhancing the model’s nonlinear fitting capability and enabling more complex expression.
The results suggest that a deeper Transformer architecture is more capable of capturing intricate patterns and dependencies within the data, which is crucial for improving error correction performance. By adding more layers, the model can perform multiple levels of abstraction and refinement, allowing it to better understand and correct errors. This is especially crucial in the context of wireless communication systems, which are characterized by complexity and variability that simple models might not adequately capture.
In essence, the findings emphasize the importance of model depth in Transformer architectures for error correction tasks. This insight can guide the design of more effective models by prioritizing the depth of the encoder layers over the dimension of $d_{k}$, leading to better performance in practical applications.
Figure 10: CBER performance across various Transformer encoder layers and the dimensions $d_{k}$.
5.2 Impact of the Sparse Masked Attention
We illustrate the impact of sparse masked attention on loss, CBER, and CBLER in Fig. 11. Specifically, sparse masked attention reduces the loss by $45\%$ and leads to faster convergence during training. Furthermore, this approach significantly improves error correction performance: the CBER shows an improvement of approximately $0.4$ dB at a bit error rate of $10^{-5}$, and the CBLER improves by about $0.5$ dB at a block error rate of $10^{-4}$.
These improvements underscore the efficacy of integrating domain-specific knowledge into the attention mechanism, demonstrating substantial gains in both efficiency and accuracy. This approach not only reduces computational complexity but also enhances the model’s ability to decode complex signals, making it a promising strategy for future advancements in error correction techniques.
Figure 11: The overall improvement in loss, CBER and CBLER with sparse masked attention.
@@ 5.3 Computational Complexity Analysis
Before delving into the computational complexity, let’s first review the unified error correction code Transformer presented in Figure 7, which incorporates $L$ layers of Transformer encoders. The principal computational complexity within each layer is found in the multi-head self-attention module, which is predominantly responsible for matrix multiplication and is defined by the hyperparameters $H,d_{k}$ and $d_{l}$. Consequently, the computational complexity of the proposed architecture is given by $O\left({L\times H\times\left({{H_{d}}\times{d_{k}}\times{d_{l}}\times N+d_{k}^{2}}\right)}\right)$, where $N=2n-k$, $d_{l}\ll N$ and $H\ll N$. Thanks to the sparsity of the parity-check matrix, $H_{d}$ is usually quite small. In our experiment, $H_{d}$ is approximately 0.17, which significantly reduces the computational complexity of the neural network.
In comparison, the approach in [22], which employs the vanilla Transformer architecture, has a complexity of $O\left({L\times H\times\left({{N^{2}}\times{d_{k}}+d_{k}^{2}}\right)}\right)$. When it comes to non-learning solutions, the computational complexities of the SCL and NMS decoding algorithms are $O\left(L_{s}\times n\times log_{2}\left(n\right)\right)+O\left(L_{s}\times\left(n-1\right)\right)+k\times O\left(2\times L_{s}\times log_{2}\left(2\times L_{s}\right)\right)$ and $I_{max}\times O\left(2\times n\times d_{\nu}+M\times\left(2\times d_{c}-1\right)\right)+I_{max}\times O\left(M\times d_{c}\right)$, respectively [34]. Here, $L_{s}$ is the list size, $I_{max}$ is the maximum number of iterations, $M$ is the size of the check matrix, $d_{v}$ is the number of variable node, and $d_{c}$ is the number of parity check node.
Compared to the approach in [22], while the proposed method excels in decoding versatility and computational complexity, it may fall short in terms of memory efficiency, power consumption, and computational resource utilization when compared to non-learning solutions. This inefficiency could limit its potential for deployment. To address these issues, future implementations could leverage techniques such as parameter sharing, low-rank factorization, and quantization, aiming to reduce complexity and enhance efficiency [35, 36].
@@ 6 Conclusions
We introduced a novel, unified code-agnostic Transformer decoding architecture, designed to integrate various linear block codes seamlessly and offer superior decoding performance. Our approach leverages the self-attention mechanism inherent in Transformer models, enabling effective handling of sequential data and capturing long-range dependencies critical for decoding tasks. Through rigorous analysis and extensive experimental validation, our results indicate that the unified Transformer decoding architecture can significantly advance the state-of-the-art in decoding performance, potentially setting new benchmarks for future research and practical applications.

### v2

@@ 5 Analysis
We investigate the impact of Transformer encoder layers on performance, focusing on the embedding dimension $d_{k}$ and the effectiveness of sparse masked attention. Additionally, we compare the computational complexity of our approach with existing methods.
5.1 Impact of Encoder Layers and Embedding Dimensions
Figure 11 displays the ABER performance across varying encoder layers $L$ and embedding dimensions $d_{k}$. Results reveal that $L$ exerts a stronger influence on performance than $d_{k}$. Increasing $L$ broadens the parameter space, enhancing nonlinear fitting and enabling complex representations.
Figure 11: ABER performance across various Transformer encoder layers and the embedding dimensions $d_{k}$.
Deeper architectures better capture intricate data patterns and dependencies, crucial for error correction. Additional layers enable multi-level abstraction and refinement, improving error interpretation and correction, especially in complex wireless communication systems where simpler models falter.
These findings highlight the primacy of model depth in Transformer-based error correction. Prioritizing $L$ over $d_{k}$ can optimize model design, enhancing practical performance.
5.2 Impact of Sparse Masked Attention
Figure 12 illustrates the effect of sparse masked attention on training loss, ABER, and ABLER. Sparse masked attention reduces training loss by 33.3% and speeds up convergence. It also significantly boosts error correction, improving ABER by 0.7 dB at a bit error rate of $10^{-5}$ and ABLER by 0.9 dB at a block error rate of $10^{-4}$.
Figure 12: The overall improvement in loss, ABER and ABLER with sparse masked attention.
These gains highlight the effectiveness of embedding domain-specific knowledge into the attention mechanism, yielding significant improvements in efficiency and accuracy. By reducing computational complexity while boosting the model’s capacity to decode complex signals, this approach emerges as a promising advancement for future error correction techniques.
@@ 5.3 Computational Complexity Analysis
Before delving into the computational complexity, let’s first review the unified error correction code Transformer presented in Figure 8, which incorporates $L$ layers of Transformer encoders. The principal computational complexity within each layer is found in the multi-head self-attention module, which is predominantly responsible for matrix multiplication and is defined by the hyperparameters $H$, $d_{k}$ and $d_{l}$. Consequently, the computational complexity of the proposed architecture is given by $O\left(L\times H\times\left(N\times d_{l}\times d_{k}\times H_{d}\right)\right)$, where $N=2n-k$, and $d_{l}\ll N$. Thanks to the sparsity of the parity-check matrix, $H_{d}$ is usually quite small. In our experiment, $H_{d}$ is approximately 0.14, which significantly reduces the computational complexity of the neural network. By contrast, the vanilla Transformer architectures in ECCT and FECCT have complexities of $O\left(L\times H\times\left(N^{2}\times d_{k}\times M_{d}+N\times d_{k}^{2}\right)\right)$ and $O\left(L\times H\times\left(N^{2}\times d_{k}+N\times d_{k}^{2}+N^{2}\times M_{d}\right)\right)$, respectively, where $M_{d}$ represents the mask density in these models.
Table 2: Comparison of trainable parameters, MACs, training time, inference time, and mask density among ECCT, FECCT, and UECCT for various codes, all with identical parameters $L=6$, $H=8$, $d_{k}=64$, $d_{l}=64$, and $d_{f}=2048$
Table 2 compares the Trainable Parameters (TPs), Multiply-Accumulate Operations (MACs), training time, inference time, and mask density for ECCT, FECCT, and UECCT. Lower values are preferable, with the best performing values highlighted in bold. The metrics for TPs and MACs are obtained using Thop [42], a neural network profiling tool. UECCT reduces parameters and MACs by 25.0% compared to both ECCT and FECCT through low-rank approximation, shared attention, and omission of Q and K projections. This results in training times 19.7% and 81.2% lower, and inference times 23.1% and 93.5% lower than ECCT and FECCT, respectively. Mask densities are influenced by design choices: FECCT has the highest density due to Tanner graph distance calculations, ECCT has a moderate density from extended bit-pair relations, and UECCT has the lowest density, 58.2% and 79.7% lower than ECCT and FECCT, achieved through the direct utilization of H and the extension of the identity matrix. For a more intuitive representation, the average values of each metric across the codewords in Table 2 are provided in Figure 13.
Figure 13: Comparison of average TPs, MACs, training time, inference time, and mask density between ECCT, FECCT and UECCT.
Compared to the approach in [24, 25], while the proposed method excels in decoding versatility and computational complexity, it may fall short in terms of memory efficiency, power consumption, and computational resource utilization when compared to non-learning solutions. This inefficiency could limit its potential for deployment. To address these issues, future implementations could leverage techniques such as parameter sharing, low-rank factorization, and quantization, aiming to reduce complexity and enhance efficiency [43, 44].
@@ 6 Conclusions
We introduced a unified, code-agnostic Transformer decoding architecture, designed to integrate various linear block codes seamlessly and offer superior decoding performance. Our approach leverages the self-attention mechanism inherent in Transformer models, enabling effective handling of sequential data and capturing long-range dependencies critical for decoding tasks. Through rigorous analysis and extensive experimental validation, our results indicate that the unified Transformer decoding architecture can significantly advance the state-of-the-art in decoding performance, potentially setting new benchmarks for future research and practical applications.